# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v2

@@ Abstract
We present masked Lagrange reconstruction, a technique that enables threshold ML-DSA (FIPS 204) with arbitrary thresholds $T$ while producing standard 3.3 KB signatures verifiable by unmodified FIPS 204 implementations. Concurrent approaches have limitations: Bienstock et al. (ePrint 2025/1163) achieve arbitrary $T$ but require honest-majority and 37–136 rounds; Celi et al. (ePrint 2026/013) achieve dishonest-majority but are limited to $T\leq 6$. Our technique addresses the barrier that Lagrange coefficients grow as $\Theta(q)$ for moderate $T$, making individual contributions too large for ML-DSA’s rejection sampling.
Unlike ECDSA threshold schemes where pairwise masks suffice for correctness, ML-DSA requires solving three additional challenges absent in prior work: (1) rejection sampling on $\|\mathbf{z}\|_{\infty}$ must still pass after masking, (2) the r0-check exposes $c\mathbf{s}_{2}$ enabling key recovery if unprotected, and (3) the resulting Irwin-Hall nonce distribution must preserve EUF-CMA security. We solve all three.
We instantiate this technique in three deployment profiles with full security proofs. Profile P1 (TEE-assisted) achieves 3-round signing with a trusted coordinator, with EUF-CMA security under Module-SIS (Theorem 4.2). Profile P2 (fully distributed) eliminates hardware trust via MPC in 8 rounds, achieving UC security against malicious adversaries corrupting up to n1n-1 parties (Theorem 6.2). Profile P3 (2PC-assisted) uses lightweight 2PC for the r0-check in 3–5 rounds, achieving UC security under a 1-of-2 CP honest assumption (Theorem 6.5) with the best empirical performance (249ms).
Our scheme requires $|S|\geq T+1$ signers and achieves success rates of 23–32%, matching single-signer ML-DSA.
@@ 1.2 Our Contribution
We present Masked Threshold ML-DSA, a threshold signature scheme for ML-DSA that supports arbitrary thresholds $T\leq N$ while producing standard-size signatures ($\approx 3.3$ KB for ML-DSA-65) identical to single-signer ML-DSA. Our scheme requires a two-honest requirement: signing sets must have size $|S|\geq T+1$ to ensure at least two honest parties for mask privacy. This addresses the threshold gap for $T\in[9,32]$ while remaining practical—configurations like $(3,5)$, $(11,15)$, and $(16,31)$ naturally satisfy $N\geq T+1$.
The scheme achieves practical success rates comparable to single-signer ML-DSA (theoretical baseline $\approx 20\%$, observed $3232\%$ due to the more concentrated Irwin-Hall nonce distribution) and is fully verifier-compatible with FIPS 204: produced signatures are accepted by unmodified FIPS 204 verifiers. While our signer-side protocol is interactive and induces an Irwin-Hall nonce distribution rather than uniform, this affects only the signing process, not verification.
@@
which correctly reconstructs the signature vector $\mathbf{z}$ and passes rejection sampling with probability matching single-signer ML-DSA (theoretical $\approx 20\%$, observed $3232\%$ for ML-DSA-65). We note that while the aggregated nonce $\mathbf{y}=\sum_{i}\mathbf{y}_{i}$ follows an Irwin-Hall distribution rather than uniform (see Definition O.5), this does not affect EUF-CMA security and only introduces a quantifiable statistical difference that we analyze in Section 3.
@@
Our implementation achieves practical signing times: approximately 1 second average for $(3,5)$-threshold and 4 seconds for $(16,31)$-threshold on commodity hardware (including retries due to rejection sampling), using NTT-accelerated polynomial arithmetic. The communication cost is approximately 12.3 KB per party per signing attempt.
@@
A fully untrusted coordinator where the r0-check is computed via an MPC subprotocol using SPDZ [DPSZ12] and edaBits [EKMOZ20]. We prove UC security against malicious adversaries corrupting up to n1n-1 parties in 8 online rounds (Theorem 6.2).
Profile P3 (2PC-Assisted):
A practical middle ground where two designated Computation Parties (CPs) jointly evaluate the r0-check via lightweight 2PC. Motivated by recent round-reduction efforts in threshold ECDSA [LLZD25], we pursue a similar goal for the lattice setting; while their Non-Interactive Multiplication technique does not directly apply to ML-DSA’s r0-check structure, we achieve analogous round improvements via 2PC. We prove UC security under the assumption that at least one CP is honest (Theorem 6.5)—weaker than TEE trust, more efficient than full MPC. Achieves the best empirical performance with 3–5 rounds and 249ms signing time.
@@ 1.4 Comparison with Prior and Concurrent Work
Scheme $T$ Size Rnd FIPS Dishon. UC
del Pino-Niot [dPN25] $\leq 8$ 3.3 KB 3 $\times$
Noise Flooding [BCDG21] $\infty$ 17 KB 3 $\times$ $\times$
Ringtail [BMMSZ25] $\infty$ 13 KB 2 $\times$ $\times$
@@
Bienstock et al. [BDEPT25] $\infty$ 3.3 KB 37–136 $\times$
Celi et al. [CDENP26] $\leq 6$ 3.3 KB 6 $\times$
Trilithium [Trilithium25] 2 3.3 KB 3 $\times$
@@
This work (P2) $\infty$ 3.3 KB 8
This work (P3) $\infty$ 3.3 KB 3–5 (1/2 CP)
Table 1: Comparison of threshold ML-DSA schemes. “Dishon.” = dishonest-majority security; “(TEE)” and “(1/2 CP)” indicate trust assumptions.
@@
Profile P2 combines arbitrary thresholds, dishonest-majority security, constant rounds (8 rounds), UC security, and FIPS 204 compatibility—a combination not achieved by prior work. Profile P3 also achieves UC security (Theorem 6.5) under the weaker 1-of-2 CP honest assumption, with the best performance (249ms, 3–5 rounds).
We contribute: (1) Masked Lagrange reconstruction, adapting pairwise-canceling masks [CGGMP20] to lattices while addressing rejection sampling, r0-check key leakage, and Irwin-Hall nonce security; (2) Three deployment profiles (P1-TEE, P2-MPC, P3-2PC) with different trust/performance trade-offs; (3) Security proofs reducing to Module-SIS/LWE, including UC security for P2 and P3; (4) A reference implementation with benchmarks from $(3,5)$ to $(16,31)$ threshold configurations.
@@ 3.2 Threat Model and Deployment Profiles
We consider a $(T,N)$-threshold setting where an adversary may corrupt up to T1T-1 parties. Table 2 summarizes our two deployment profiles.
@@
Property Profile P1 Profile P2 Profile P3
@@
Online rounds 3 8 3–5
Latency Low High Lowest
@@
A fully distributed protocol where no single party learns $c\mathbf{s}_{2}$. The r0-check is computed via an MPC subprotocol using SPDZ [DPSZ12] and edaBits [EKMOZ20] (see Section 6). This provides stronger distributed trust at the cost of additional rounds (8 online rounds). We prove UC security against malicious adversaries corrupting up to n1n-1 parties in Theorem 6.2.
Profile P3 (2PC-Assisted).
A hybrid approach that designates two Computation Parties (CPs) to jointly evaluate the r0-check via lightweight 2PC. Each CP receives additive shares of $c\mathbf{s}_{2}$; neither learns the complete value unless they collude. This achieves the best empirical performance (see Table 5) while requiring only that at least one of the two CPs is honest—a weaker assumption than TEE trust and more efficient than full MPC. The 2PC uses standard garbled circuits [Yao86] or oblivious transfer-based protocols for the infinity norm comparison. Profile P3 is ideal for deployments where two semi-trusted servers can be provisioned (e.g., multi-cloud setups), offering a practical middle ground between P1’s hardware trust and P2’s computational overhead.
@@ 3.7 Nonce Distribution and Security
The aggregated nonce $\mathbf{y}=\sum_{i\in S}\mathbf{y}_{i}$ follows an Irwin-Hall distribution concentrated around zero. This preserves EUF-CMA security: using smooth Rényi divergence bounds from [Raccoon2024], we verify $R^{\epsilon}_{2}\leq 1.003$ for $|S|\leq 33$ signers (Appendix D), yielding negligible security loss ($2^{-40}$ even for $2^{64}$ queries).
@@ 3.8 Abort Probability
The per-attempt success probability matches single-signer ML-DSA ($\approx 20\%$; empirically $\%3232\%$ due to Irwin-Hall concentration). Without masking, success degrades as $(0.2)^{T}$—for $T=16$, this is $6.6\times 10^{-12}$. Our masked approach achieves constant success rate, a speedup of $10^{5}$$10^{21}\times$ (Appendix F).
@@ 3.9 The r0-Check Subprotocol
The combiner can compute this as follows:
  • $c\mathbf{s}_{2}$ is obtained from the masked reconstruction above
  • $\mathbf{A}\mathbf{z}=\mathbf{A}(\mathbf{y}+c\mathbf{s}_{1})=\mathbf{w}+c\mathbf{A}\mathbf{s}_{1}$, so $c\mathbf{A}\mathbf{s}_{1}=\mathbf{A}\mathbf{z}-\mathbf{w}$
  • Thus $c\mathbf{t}_{0}=(\mathbf{A}\mathbf{z}-\mathbf{w})+c\mathbf{s}_{2}-c\mathbf{t}_{1}\cdot 2^{d}$
@@
We address this via three deployment profiles (Table 2):
  • P1 (TEE): $c\mathbf{s}_{2}$ computed inside enclave, never exposed
  • P2 (MPC): Parties jointly evaluate r0-check without reconstructing $c\mathbf{s}_{2}$
  • P3 (2PC): Two CPs each hold additive share; neither learns complete value
See Section 6 for P2/P3 details.
@@ 4.4 Nonce Distribution Security
Let $\Pi$ be our threshold ML-DSA with Irwin-Hall nonces. Then: $\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq R_{2}^{\epsilon}\cdot\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi^{*}}(\mathcal{A})+q_{S}\cdot\epsilon$ where $R_{2}^{\epsilon}\leq 1+n^{4}\beta^{2}/(4\gamma_{1}^{2})$ for $|S|=n$ signers. Concretely for ML-DSA-65: $R_{2}^{\epsilon}\leq 1.003$ for $|S|\leq 17$, $R_{2}^{\epsilon}\leq 1.05$ for $|S|\leq 33$, and $\epsilon<10^{-30}$.
For $q_{S}=2^{64}$ signing queries, the multiplicative loss remains small ($\leq 1.05$ for practical thresholds) and additive loss is negligible ($2^{-36}$). Detailed calculations appear in Appendix D.
@@ 4.7 Security Across Deployment Profiles
All three deployment profiles (P1, P2, P3) achieve identical EUF-CMA security bounds, as the final signature structure $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ is independent of the profile used. Additionally, all three profiles achieve UC security under their respective trust assumptions.
@@ 4.7.1 Profile P1 UC Security
Theorem 4.10 assumes the TEE is ideal—it correctly executes the specified computation and reveals nothing beyond the output. This is a hardware trust assumption, not a cryptographic one. In practice, TEE security depends on the specific implementation (Intel SGX, ARM TrustZone, etc.) and is subject to side-channel attacks. Profiles P2 and P3 eliminate this assumption via cryptographic protocols.

### v3

@@ Abstract
[…]
Our technique overcomes the barrier that Lagrange coefficients grow as $\Theta(q)$ for moderate $T$, making individual contributions too large for rejection sampling. We introduce pairwise-canceling masks that hide these coefficients while preserving correctness. Unlike ECDSA threshold schemes, ML-DSA presents three additional challenges: the $\|\mathbf{z}\|_{\infty}$ check must pass after masking, the r0-check must not leak $c\mathbf{s}_{2}$, and the resulting Irwin-Hall nonce distribution must preserve EUF-CMA security. We solve all three with complete UC proofs.
We instantiate this technique in three deployment profiles. Profile P1 uses a TEE coordinator (3 rounds, 6ms). Profile P2 eliminates hardware trust via MPC (5 rounds, 22ms for small thresholds, up to 194ms for $T=16$). Profile P3+ uses lightweight 2PC with semi-asynchronous signing where signers precompute nonces offline and respond within a time window (2 logical rounds, 22ms), similar to FROST.
Our Rust implementation scales from 2-of-3 to 32-of-45 thresholds. P1 and P3+ achieve sub-100ms latency for all configurations; P2 trades higher latency for eliminating hardware trust. Success rates of $\approx$21–45% are comparable to single-signer ML-DSA’s $\approx$20–25%.
@@ 1.2 Our Contribution
We present Masked Threshold ML-DSA, a threshold signature scheme for ML-DSA that supports arbitrary thresholds $T\leq N$ while producing standard-size signatures ($\approx 3.3$ KB for ML-DSA-65) identical to single-signer ML-DSA. Our scheme requires a two-honest requirement: signing sets must have size $|S|\geq T+1$ to ensure at least two honest parties for mask privacy. This addresses the threshold gap for $T\in[9,32]$ while remaining practical; configurations like $(3,5)$, $(11,15)$, and $(16,21)$ naturally satisfy $N\geq T+1$.
The scheme achieves practical success rates comparable to single-signer ML-DSA (theoretical baseline $\approx 20\%$, observed $\approx$21–45% due to the more concentrated Irwin-Hall nonce distribution) and is fully verifier-compatible with FIPS 204: produced signatures are accepted by unmodified FIPS 204 verifiers. While our signer-side protocol is interactive and induces an Irwin-Hall nonce distribution rather than uniform, this affects only the signing process, not verification.
@@
which correctly reconstructs the signature vector $\mathbf{z}$ and passes rejection sampling with probability comparable to single-signer ML-DSA (theoretical $\approx 20\%$, observed $\approx$21–45% for ML-DSA-65). We note that while the aggregated nonce $\mathbf{y}=\sum_{i}\mathbf{y}_{i}$ follows an Irwin-Hall distribution rather than uniform (see Definition Q.5), this does not affect EUF-CMA security and only introduces a quantifiable statistical difference that we analyze in Section 3.
@@
Our optimized Rust implementation achieves practical signing times: 5.8ms (P1-TEE), 21.5ms (P2-MPC), and 22.1ms (P3+-Async) for $(3,5)$-threshold on commodity hardware (including retries). P1 and P3+ remain under 100ms for all configurations up to $(32,45)$; P2 trades higher latency ($\sim$130–194ms for $T\geq 16$) for eliminating hardware trust. The communication cost is approximately 12.3 KB per party per signing attempt.
@@
A fully untrusted coordinator where the r0-check is computed via an MPC subprotocol using SPDZ [DPSZ12] and edaBits [EKMOZ20]. Using combiner-mediated commit-then-open and constant-depth comparison circuits, we achieve 5 online rounds (reduced from 8). We prove UC security against malicious adversaries corrupting up to n1n-1 parties (Theorem 6.2).
Profile P3+ (Semi-Async 2PC):
Designed for human-in-the-loop authorization. Two designated Computation Parties (CPs) jointly evaluate the r0-check via 2PC, but signers participate semi-asynchronously: they precompute nonces offline and respond within a time window (e.g., 5 minutes) rather than synchronizing for multiple rounds. This achieves 2 logical rounds (1 signer step + 1 server round). We prove UC security under the assumption that at least one CP is honest (Theorem 6.5). The “Async” column in Table 1 distinguishes our work from concurrent approaches, which require synchronous multi-round signer participation.
@@ 1.4 Comparison with Prior and Concurrent Work
Scheme $T$ Size Rnd FIPS Dishon. UC Async
del Pino-Niot [dPN25] $\leq 8$ 3.3 KB 3 $\times$ $\times$
Noise Flooding [BCDG21] $\infty$ 17 KB 3 $\times$ $\times$ $\times$
Ringtail [BMMSZ25] $\infty$ 13 KB 2 $\times$ $\times$ $\times$
@@
Bienstock et al. [BDEPT25] $\infty$ 3.3 KB 37–136 $\times$ $\times$
Celi et al. [CDENP26] $\leq 6$ 3.3 KB 6 $\times$ $\times$
Trilithium [Trilithium25] 2 3.3 KB 3 $\times$ $\times$
@@
This work (P2) $\infty$ 3.3 KB 5 $\times$
This work (P3+) $\infty$ 3.3 KB 2 (1/2 CP)
Table 1: Comparison of threshold ML-DSA schemes. ✓= supported, $\times$ = not supported. “Dishon.” = dishonest-majority security (tolerates n1n-1 corruptions); Bienstock et al. require honest majority. “(TEE)” and “(1/2 CP)” indicate additional trust assumptions beyond threshold. “Async” = semi-asynchronous signer participation. P3+ requires 1 signer step + 1 server round = 2 logical rounds. P1 UC security assumes ideal TEE (Theorem 4.10).
@@
Profile P2 combines arbitrary thresholds, dishonest-majority security, constant rounds (5 rounds), UC security, and FIPS 204 compatibility, a combination not achieved by prior work. Profile P3+ achieves UC security (Theorem 6.5) under the 1-of-2 CP honest assumption, with semi-asynchronous signer participation (2 logical rounds).
We contribute: (1) Masked Lagrange reconstruction, adapting pairwise-canceling masks [CGGMP20] to lattices while addressing rejection sampling, r0-check key leakage, and Irwin-Hall nonce security; (2) Three deployment profiles (P1-TEE, P2-MPC, P3+-2PC) with different trust/performance trade-offs, including the novel semi-async model for P3+; (3) Security proofs reducing to Module-SIS/LWE, including complete UC security proofs for P2 and P3+; (4) A reference implementation with benchmarks from $(2,3)$ to $(32,45)$ threshold configurations.
@@ 3.2 Threat Model and Deployment Profiles
We consider a $(T,N)$-threshold setting where an adversary may corrupt up to T1T-1 parties. Table 2 summarizes our three deployment profiles.
@@
Property Profile P1 Profile P2 Profile P3+
@@
Online rounds 3 5 2
Signer sync Multi-round Multi-round Semi-async
Latency Low Medium Lowest
@@
P3+ = 1 signer step (within time window) + 1 server round (2PC). Signers precompute nonces offline.
@@
A fully distributed protocol where no single party learns $c\mathbf{s}_{2}$. The r0-check is computed via an MPC subprotocol using SPDZ [DPSZ12] and edaBits [EKMOZ20] (see Section 6). Using combiner-mediated commit-then-open and constant-depth comparison circuits, we achieve 5 online rounds (optimized from 8). We prove UC security against malicious adversaries corrupting up to n1n-1 parties in Theorem 6.2.
Profile P3+ (Semi-Async 2PC).
Designed for human-in-the-loop authorization scenarios. Two designated Computation Parties (CPs) jointly evaluate the r0-check via lightweight 2PC. Each CP receives additive shares of $c\mathbf{s}_{2}$; neither learns the complete value unless they collude.
The key innovation is semi-asynchronous signer participation: signers precompute nonces $(\mathbf{y}_{i},\mathbf{w}_{i},\mathsf{Com}_{i})$ offline at any time, then respond within a bounded time window (e.g., 5 minutes) after receiving the challenge. A client daemon handles the cryptographic operations transparently, completing the signing process in under 1 second.
This achieves 2 logical rounds (1 signer step + 1 server round). The 2PC uses pre-garbled circuits for reduced online latency ($\sim$15ms). Profile P3+ is suited for deployments requiring human authorization where signers cannot coordinate for multiple synchronous rounds.
@@ 3.7 Nonce Distribution and Security
The aggregated nonce $\mathbf{y}=\sum_{i\in S}\mathbf{y}_{i}$ follows an Irwin-Hall distribution concentrated around zero. This preserves EUF-CMA security: using smooth Rényi divergence bounds from [Raccoon2024], we verify $R^{\epsilon}_{2}\leq 1.003$ for $|S|\leq 17$ and $R^{\epsilon}_{2}\leq 1.05$ for $|S|\leq 33$ signers (Appendix D), yielding negligible security loss ($2^{-40}$ even for $2^{64}$ queries).
@@ 3.8 Abort Probability
The per-attempt success probability is comparable to single-signer ML-DSA ($\approx$20%; empirically $\approx$21–45% due to Irwin-Hall concentration and configuration-dependent variance). Without masking, success degrades as $(0.2)^{T}$—for $T=16$, this is $6.6\times 10^{-12}$. Our masked approach achieves constant success rate, a speedup of $10^{5}$$10^{21}\times$ (Appendix F).
@@ 3.9 The r0-Check Subprotocol
The combiner can compute this as follows. First, $c\mathbf{s}_{2}$ is obtained from the masked reconstruction above. Second, since $\mathbf{A}\mathbf{z}=\mathbf{A}(\mathbf{y}+c\mathbf{s}_{1})=\mathbf{w}+c\mathbf{A}\mathbf{s}_{1}$, we have $c\mathbf{A}\mathbf{s}_{1}=\mathbf{A}\mathbf{z}-\mathbf{w}$. Combining these, $c\mathbf{t}_{0}=(\mathbf{A}\mathbf{z}-\mathbf{w})+c\mathbf{s}_{2}-c\mathbf{t}_{1}\cdot 2^{d}$.
@@
We address this via three deployment profiles (Table 2). In Profile P1 (TEE), $c\mathbf{s}_{2}$ is computed inside the enclave and never exposed. In Profile P2 (MPC), parties jointly evaluate the r0-check without reconstructing $c\mathbf{s}_{2}$. In Profile P3+ (2PC), two CPs each hold an additive share, so neither learns the complete value. See Section 6 for P2/P3 details.
@@ 4.4 Nonce Distribution Security
Let $\Pi$ be our threshold ML-DSA with Irwin-Hall nonces. Then: $\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\sqrt{\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi^{*}}(\mathcal{A})+\epsilon}\cdot R_{2}^{\epsilon}+q_{S}\cdot\epsilon$ where $R_{2}^{\epsilon}\leq 1+n^{4}\beta^{2}/(4\gamma_{1}^{2})$ for $|S|=n$ signers. Concretely for ML-DSA-65: $R_{2}^{\epsilon}\leq 1.003$ for $|S|\leq 17$, $R_{2}^{\epsilon}\leq 1.05$ for $|S|\leq 33$, and $\epsilon<10^{-30}$.
The square root arises from the generic Rényi divergence probability bound; specialized reductions for Fiat-Shamir with Aborts [Raccoon2024] can achieve tighter (linear) bounds. With ML-DSA-65’s 192-bit security ($\mathsf{Adv}_{\Pi^{*}}\leq 2^{-192}$), the bound gives $\mathsf{Adv}_{\Pi}\leq 1.05\cdot 2^{-96}+2^{-36}<2^{-35}$, maintaining adequate security margin. Detailed calculations appear in Appendix D.
@@ 4.7 Security Across Deployment Profiles
All three deployment profiles (P1, P2, P3+) achieve identical EUF-CMA security bounds, as the final signature structure $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ is independent of the profile used. Additionally, all three profiles achieve UC security under their respective trust assumptions.
@@ 4.7.1 Profile P1 UC Security
Theorem 4.10 assumes the TEE is ideal—it correctly executes the specified computation and reveals nothing beyond the output. This is a hardware trust assumption, not a cryptographic one. In practice, TEE security depends on the specific implementation (Intel SGX, ARM TrustZone, etc.) and is subject to side-channel attacks. Profiles P2 and P3+ eliminate this assumption via cryptographic protocols.
4.7.2 Profile P3+ Semi-Async Security Model
Profile P3+ introduces a semi-asynchronous execution model where signers participate within a bounded time window rather than synchronizing for multiple rounds. This requires extending the UC framework to handle timing.
Definition 4.13 (Semi-Async Timing Model).
The semi-async model defines the following timing structure:
  • Pre-signing window $t_{0},t_{1}]$: Signers may precompute nonces $(y_{i},w_{i},\mathsf{Com}_{i})$ at any time
  • Challenge broadcast $t_{2}$: Combiner derives challenge $c$ from aggregated commitments
  • Response window $t_{2},t_{2}+\Delta]$: Signers respond within window $\Delta$ (e.g., 5 minutes)
  • 2PC execution $t_{3},t_{4}]$: Synchronous between CPs only
The adversary cannot control timing within honest parties’ windows but may observe message arrival times.
Theorem 4.14 (Profile P3+ UC Security).
Under the semi-async timing model, Protocol P3+ UC-realizes $\mathcal{F}_{\mathsf{ThreshSig}}$ against static adversaries corrupting up to T1T-1 signers and at most one CP, assuming:
  1. 1.
    Garbled circuits are secure (Yao’s protocol)
  2. 2.
    Oblivious transfer is UC-secure
  3. 3.
    PRF is secure
  4. 4.
    At least one CP remains honest
Security bound: $\epsilon\leq\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{OT}}+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
The proof appears in Appendix P. The key insight is that pre-signing security follows from commitment hiding (signers commit to nonces before seeing the challenge), and the response window does not introduce new attack vectors because:
  1. 1.
    Replay prevention: Each signing attempt uses a unique session ID
  2. 2.
    Commitment binding: Signers cannot change their nonce after commitment
  3. 3.
    Window expiry: Late responses are rejected, preventing adaptive attacks